OpenStudy (anonymous):

Equivalent rate~ $33,000/1year =$2750/1month How did I get this?

OpenStudy (anonymous):

1year = 12 months

jimthompson5910 (jim_thompson5910):

1 year = 12 months So $33,000 /1 year = $33,000/12 months Now divide each component by 12 to get 33,000/12 = 2750 and 12/12 = 1 So $33,000/12 months = $2750/ 1 month This means that $33,000 /1 year = $2750/ 1 month

OpenStudy (anonymous):